# Flexible-length Text Infilling for Discrete Diffusion Models

### v1

@@ Abstract
Discrete diffusion models are a new class of text generators that offer advantages such as bidirectional context use, parallelizable generation, and flexible prompting compared to autoregressive models. However, a critical limitation of discrete diffusion models is their inability to perform flexible-length or flexible-position text infilling without access to ground-truth positional data. We introduce DDOT (Discrete Diffusion with Optimal Transport Position Coupling), the first discrete diffusion model to overcome this challenge. DDOT jointly denoises token values and token positions, employing a novel sample-level Optimal Transport (OT) coupling. This coupling preserves relative token ordering while dynamically adjusting the positions and length of infilled segments, a capability previously missing in text diffusion. Our method is orthogonal to existing discrete text diffusion methods and is compatible with various pretrained text denoisers. Extensive experiments on text infilling benchmarks such as One-Billion-Word and Yelp demonstrate that DDOT outperforms naive diffusion baselines. Furthermore, DDOT achieves performance on par with state-of-the-art non-autoregressive models and enables significant improvements in training efficiency and flexibility.
@@
Andrew Zhang, Anushka Sivakumar, Chiawei Tang, Chris Thomas Virginia Tech, Virginia, USA azhang42vt.edu
@@ 1 Introduction
While autoregressive models are effective, they also involve highly sequential sampling, cannot take bidirectional context, and lead to constrained architectures by requiring a decoder mask. In contrast, discrete diffusion models Lou et al. ; Sahoo et al. (2024) can parallelize generation by denoising multiple tokens simultaneously, use bidirectional information, do not need a decoder mask, and, notably, allow for more controllable generation. Previous work Gat et al. ; Shi et al. (2024) has demonstrated that discrete diffusion models are capable of arbitrary prompting locations, whereas autoregressive models are only capable of left-to-right text completion. However, this advantage has a significant limitation: existing diffusion models cannot alter the distances between these prompt tokens. Consequently, existing text diffusion models cannot generate the ground-truth sample without access to the oracle positions of the prompt and infilled text.
Figure 1: Our novel diffusion across token positions enables dynamic token movement for infilling. Unlike prior methods, DDOT learns to move mask tokens to appropriate locations, such as to the right of "brown," even if initially unmasked there. The OT coupling (colored lines) significantly simplifies this learning by drastically reducing possible permutations.
We solve this issue by enabling discrete diffusion models to learn where to move tokens. Specifically, we design a diffusion process that operates across token positions, allowing the model to vary the position and length of infilled spans. Furthermore, given the importance of token positioning in preserving semantic meaning He et al. , we incorporate sample-level OT (optimal transport) coupling to maintain relative token ordering throughout the diffusion process. Even minor positional changes can dramatically alter meaning, as seen in phrases like "The child’s green coat" and "The green child’s coat". DDOT’s OT coupling preserves this relative ordering throughout the diffusion process while supporting flexible-length text infilling. Our OT coupling prevents such swaps and drastically improves DDOT’s downstream performance across all studied benchmarks and metrics. Extensive experiments show that DDOT outperforms naive diffusion baselines and achieves on par performance with state-of-the-art non-autoregressive (NAR) models.
@@
  • We provide extensive experiments on DDOT that show it achieves outperforms diffusion baselines and achieves performance on par with state-of-the-art NAR models.
  • We provide detailed ablations and visualizations that verify DDOT’s effectiveness for adjusting the position and length of infilled text spans and provide insights into our novel sample-level OT coupling. The OT coupling significantly outperforms naive diffusion across all tested benchmarks and metrics.
Figure 2: DDOT learns to vary infilled span lengths and positions unlike prior fixed-position diffusion methods. (Left) We compute two separate intra-set OT couplings within the prompt positions and the response positions. This constraint drastically simplifies possible permutations. Right Given a timestep $t$, we predict the token and position.
@@ 2 Related Work
2.1 Lexicographically Constrained Generation
Constrained text generation has been explored through a variety of approaches, including AR and NAR methods Zhang et al. (2020); Iso (2024); He (2021); Stern et al. (2019); Lu et al. (2022). POINTER Zhang et al. (2020) enables flexible token generation through iterative insertion, though it still depends on sequential token prediction and multiple forward passes. AutoTemplate Iso (2024) approaches constrained text generation by simply feeding the prompt tokens into an encoder-decoder style model. CBART He (2021) extends the POINTER architecture by moving to an encoder-decoder model. Autoregressive methods, while effective for their specific use cases, inherit fundamental limitations: they require sequential generation that scales linearly with sequence length, and their causal attention masks prevent full utilization of bidirectional context during generation. Most critically, for text infilling tasks, these approaches struggle to simultaneously consider both past and future context when generating intermediate content Cao et al. (2023).
@@ 2.2 Discrete Diffusion Models
Discrete diffusion models offer an innovative approach to text generation, addressing key limitations of autoregressive methods Lou et al. ; Ren et al. (2024); Gong et al. (2024); Sahoo et al. (2024). These models denoise corrupted text sequences, enabling parallel updates of multiple tokens rather than the token-by-token process of autoregressive methods, reducing the number of forward passes. Additionally, their bidirectional nature allows them to leverage both past and future context in contrast to causal masking that constrains autoregressive models. Early frameworks like D3PM Austin et al. adapted continuous diffusion to discrete tokens using Markovian corruption, while subsequent advances such as multinomial diffusion Hoogeboom et al. (2021) improved noise schedules and training efficiency.
Recent work on score-based discrete diffusion has further advanced the field by providing analytical solutions for the denoising process. Instead of directly modeling transition probabilities SEDD Lou et al. uses a score-based approach that learn the gradient of the log probability, which proves particularly effective for handling high-dimensional categorical data like text.
@@ 2.3 Optimal Transport Coupling for Flexible Generation
OT Villani et al. (2009) coupling has been well studied in image generation through theoretical foundations such as Rectified Flow, Liu et al. , which showed how continuous normalizing flows can be understood as OT paths. This framework demonstrated that probability mass can be efficiently transported through straight trajectories while preserving structural relationships, providing insights into how sequential data can be manipulated while maintaining order.
Building on these insights, Tong et al. showed how OT can be incorporated into diffusion models to accelerate generation. By defining appropriate transport costs between positions, OT coupling ensures that dimensions within a sample move in straight, predictable paths.
Although previous works incorporated OT into diffusion models with a focus on generation speed, our work leverages these properties to enable flexible-length text infilling. By coupling the discrete token values with continuous position variables through OT, we allow the model to optimize both token content and positioning simultaneously. This approach maintains the parallel generation and bidirectional context advantages of discrete diffusion while adding the ability to dynamically adjust sequence lengths. The OT coupling effectively serves as a bridge between the discrete token space and continuous position space, enabling natural language generation that respects both local token relationships and global sequence structure.
@@ 3 Background: Masked Text Diffusion
Discrete diffusion adapts the continuous nature of diffusion processes to text. The simplest and most performant version of discrete diffusion is masked diffusion. Rather than adding gaussian noise to continuous values such as pixels, masked text diffusion assigns a certain probability of masking tokens throughout the forward diffusion process. For the sake of this paper, masked diffusion can be seen as a masked language model (like BERT Devlin et al. (2019)) that works at gradually decreasing masking ratios to generate text. Specifically, our text diffusion process follows Score Entropy Discrete Diffusion (SEDD) Lou et al. , modeling a score function across a support of $N$ states or token values. The forward diffusion process is given by a continuous time Markov chain (CTMC):
$\frac{dp_{t}}{dt}=Q_{t}p_{t}\quad p_{0}\approx p_{\mathrm{data}}$ (1)
$Q_{t}\in\mathbb{R}^{n\times n}$ denotes a transition matrix where the columns give the probability of transitioning from one discrete state (or token) to another. In masking diffusion models, the diagonal is $-1$, the last row is $1$, and all other elements are 00, moving probability mass from any token value to the last token in the vocabulary, the mask token. To extend the process to sequences rather than individual tokens, SEDD applies discrete diffusion independently to all tokens in a sequence.
SEDD reverses the textual diffusion equation by modeling the score $s_{\theta}(x)_{y}$ of transitioning from token $x$ to token $y$ with a score entropy loss:
${\scriptstyle\mathcal{L}_{\text{tok}}=\mathbb{E}_{(x,x_{0})}\Big{[}\sum_{y\neq x}w_{xy}\Big{(}s_{\theta}(x)_{y}-\frac{p(y|x_{0})}{p(x|x_{0})}\log s_{\theta}(x)_{y}\Big{)}\Big{]}}$ (2)
where $w_{xy}$ weighs the importance of different states. Finally, to simulate the reverse diffusion process, we either take Euler steps or use an analytical denoiser Lou et al. .
@@ 4 Approach
Previous discrete diffusion models must know the ground truth positions of mask tokens to infill a sequence, as the positions of tokens are fixed throughout the entire diffusion process. To address this limitation, we introduce DDOT, a diffusion framework that jointly denoises discrete token values and continuous token positions, allowing for flexible infilling spans.
@@ 4.1 Continuous Position Diffusion
DDOT gradually transforms token positions from a simple initial distribution to their ground truth permutation. To achieve this, we denoise $l$ token positions $z_{t}\in[-1,1]^{l}$ in the continuous domain. The initial (limiting) distribution of positions is sampled from a uniform distribution, $z_{T}\sim\mathcal{U}(-1,1)^{L}$, where $L$ represents the maximum sequence length. The ground truth positions are defined as:
$z_{0}=linspace\left(-\frac{l}{L},\frac{l}{L},l\right),$ (3)
where $z_{0}$ is evenly spaced and scaled to match the length of the true sequence. This setup ensures that the position diffusion process captures the gradual transition from a simple prior to a structured output aligned with the token ordering and length. Similar to fixed position models, we scale by $l$ to provide information on the abosolute distances between tokens.
Our preliminary studies reveal that naively performing diffusion on the token positions leads to poor performance. We observe that adding noise to token positions destroys the relative ordering of text and adds a combinatorially exploding number of permutations, problems not present in absolute-valued continuous domains such as image generation Ho et al. (2020); Karras et al. (2022). Since the relative ordering of text can drastically change its meaning and consequently abruptly destroy the signal across timesteps, it is crucial to fix token ordering across time.
Unlike traditional diffusion processes that aim to completely destroy the data signal at the final timestep $T$, we provide the ordering of the prompt tokens at $T$ to the model. To maintain the prompt token ordering, we introduce a sample level OT coupling between the initial $z_{0}$ and final $z_{T}$ token positions of the diffusion process. Additionally, we do not add noise to the positions and instead use linear interpolation between $z_{0}$ and $z_{T}$.
Following Liu et al. ; Albergo and Vanden-Eijnden , we model the forward position diffusion process with an ordinary differential equation (ODE)
$dZ_{t}=v(Z_{t},t)dt$ (4)
where $v(z_{t},t)$ is the velocity (rate of change) of the positions at time $t$. We partition $x$ and $z$ into prompt ($x^{p}$, $z^{p}$) and infilled response ($x^{r}$, $z^{r}$) subsets. The limiting positions $z_{T}$ are determined as the solution to the linear sum assignment problem, minimizing OT cost after scaling $z_{0}$in $[-1,1]$:
$z^{*}_{T}=\pi(\frac{z_{0}L}{l},z_{T}),\quad\pi=\arg\min_{\pi^{\prime}\in S_{n}}\sum_{i=1}^{n}C_{i,\pi^{\prime}(i)}$ (5)
where $C_{i,j}$ is the Euclidean distance between positions and $S_{n}$ is all possible permutations of length $n$. For prompt tokens, the optimal transport is balanced, so $|z^{p}_{T}|=|z^{p}_{0}|=l^{p}$ However, the number of infilled tokens in the limiting distribution is almost always larger than that of the ground truth $|z^{r}_{T}|=L-l^{p}\geq l-l^{p}=|z^{r}_{0}|$ To reconcile the different sizes of $z_{0}$ and $z_{T}$, we treat the remaining unassigned $L-l$ positions as pad tokens. Specifically, we set the ground truth tokens to the pad token and the ground truth positions to stationary paths before scaling: $z^{pad}_{0}=z^{pad}_{T}*\frac{l}{L}$. Importantly, these stationary paths maintain that the coupling between $z_{0}$ and $z_{T}$ achieves the optimal OT cost.
This approach establishes a coupling within prompt and infilled tokens, ensuring that tokens in each subset maintain non-crossing paths, a key property of OT Tong et al. ; Villani et al. (2009). While intra-set paths remain ordered to minimize transport cost, DDOT allows inter-set crossings (i.e., between prompt and infilled tokens), enabling flexible token positioning while preserving the relative order of prompt tokens. Path visualizations are available in A.4.
Beyond prompt tokens, DDOT ensures that relative token order within all sets, prompt and infilled response, aligns with their ground truth at any timestep. This guarantees smooth transitions and avoids disruptions caused by abrupt positional swaps. Using OT, DDOT simplifies position manipulation, reducing unnecessary degrees of freedom.
OT is traditionally computationally expensive, requiring iterative calculations on the CPU. However, our OT coupling is computationally efficient, addressing the large-scale data demands of generative text modeling. Previous methods approximate dataset-level OT through per-batch computations Tong et al. , typically relying on algorithms like the Sinkhorn method Cuturi (2013) and operating in high-dimensional spaces (e.g., image latents). In contrast, our OT coupling operates at the scalar level and limited to $L$ elements, corresponding to the model’s context size—a scale significantly smaller than the number of samples in a dataset. For prompt tokens, the OT coupling is balanced and can therefore be efficiently computed by simply sorting $z_{0}$ and $z_{T}$. This efficiency enables us to calculate exact sample-level OT, unlike the approximations required in previous dataset-level approaches.
For training, we randomly sample a timestep $t$ and retrieve the corresponding positions in a simulation-free manner with linear interpolation $z_{t}=(1-t)z_{0}+tz_{T}$. The training objective is a weighted mean-squared error loss:

### v2

@@ Abstract
Discrete diffusion models are a new class of text generation models that offer advantages such as bidirectional context, parallelizable generation, and flexible prompting compared to autoregressive models. However, a critical limitation has been the inability to perform flexible-length or flexible-position text infilling without access to ground-truth positional data. We introduce DDOT [^1] Project page: https://andrew-zhang.github.io/ddot-page (Discrete Diffusion with Optimal Transport Position Coupling), a discrete diffusion model that overcomes this limitation by jointly denoising token values and token positions using a novel sample-level optimal transport coupling. This coupling preserves relative token ordering while dynamically adjusting the positions and lengths of infilled segments. DDOT is orthogonal to existing discrete text diffusion methods and is compatible with various pretrained text denoisers. On text-infilling benchmarks such as One-Billion-Word and Yelp, DDOT outperforms naive diffusion baselines and achieves performance on par with state-of-the-art non-autoregressive models, while improving training efficiency and prompting flexibility.
@@
Andrew Zhang    Anushka Sivakumar    Chiawei Tang    Chris Thomas Department of Computer Science Virginia Tech {azhang42,anushkas01,cwtang,christhomas}@vt.edu
@@ 1 Introduction
While autoregressive (AR) models are effective, they also involve highly sequential sampling, cannot use bidirectional context, and constrain architectures by requiring a decoder mask. In contrast, discrete diffusion models Lou et al. (2024); Sahoo et al. (2024) can parallelize generation by denoising multiple tokens simultaneously, use bidirectional context, do not need a decoder mask, and allow for more controllable generation. Previous work Gat et al. (2024); Shi et al. (2024) has demonstrated that discrete diffusion models can handle prompts at arbitrary locations, whereas autoregressive models are only capable of left-to-right text completion. However, this advantage has a significant limitation: existing diffusion models cannot alter the distances between these prompt tokens. Consequently, existing text diffusion models cannot generate the ground-truth sample without access to the oracle positions of the prompt and infilled text.
Figure 1: Our diffusion across token positions enables dynamic token movement for infilling. Unlike prior methods, DDOT learns to move masked tokens to appropriate locations, such as to the right of “brown,” even if that position was not initially masked. The OT coupling (colored lines) simplifies this learning by drastically reducing the set of possible permutations.
We solve this issue by enabling discrete diffusion models to learn where to move tokens. Specifically, we design a diffusion process that operates across token positions (in addition to token values), allowing the model to vary the positions and lengths of infilled spans (Figure 1). Furthermore, given the importance of token positioning in preserving semantic meaning He et al. (2020), we incorporate sample-level OT coupling Tong et al. (2024) to maintain relative token ordering throughout the diffusion process. Even minor positional changes can dramatically alter meaning, as seen in phrases like “The child’s green coat” and “The green child’s coat.” DDOT’s OT coupling preserves this relative ordering throughout the diffusion process while supporting flexible-length text infilling. Our OT coupling prevents such swaps and drastically improves DDOT’s training efficiency and downstream performance across all studied benchmarks and metrics. Extensive experiments show that DDOT outperforms naive diffusion baselines and achieves performance on par with state-of-the-art non-autoregressive (NAR) models.
@@
  • We provide extensive experiments on DDOT that show it outperforms diffusion baselines and achieves performance on par with state-of-the-art NAR models.
  • We provide detailed ablations and visualizations that verify DDOT’s effectiveness in adjusting the positions and lengths of infilled text spans and provide insights into our novel sample-level OT coupling. The OT coupling significantly outperforms naive diffusion baselines across all tested benchmarks and metrics.
Figure 2: DDOT learns to vary infilled span lengths and positions, unlike prior fixed-position diffusion methods. (Left) We compute two separate intra-set OT couplings within the prompt positions and the response positions, which drastically simplifies the set of possible permutations. (Right) Given a time step $t$, we predict the token and position.
@@ 2 Related Work
2.1 Lexically Constrained Generation
Constrained text generation has been explored through a variety of approaches, including AR and NAR methods Zhang et al. (2020); Iso (2024); He (2021); Stern et al. (2019); Lu et al. (2022). POINTER Zhang et al. (2020) enables flexible token generation through iterative insertion, though it still depends on sequential token prediction and multiple forward passes. AutoTemplate Iso (2024) approaches constrained text generation by feeding the prompt tokens into an encoder–decoder-style model. CBART He (2021) extends the POINTER architecture by moving to an encoder–decoder model. Autoregressive methods, while effective for their specific use cases, inherit fundamental limitations: they require sequential generation that scales linearly with sequence length, and their causal attention masks prevent the full use of bidirectional context during generation. Most critically, for text infilling tasks, these approaches struggle to consider both past and future contexts simultaneously when generating intermediate content Cao et al. (2023). Furthermore, to allow for flexible positions, autoregressive methods often regenerate the entire sequence instead of only inserting tokens, leading to wasted compute Iso (2024).
@@ 2.2 Discrete Diffusion Models
Discrete diffusion models offer an innovative approach to text generation, addressing key limitations of autoregressive methods Lou et al. (2024); Ren et al. (2024); Gong et al. (2024); Sahoo et al. (2024). These models denoise corrupted text sequences, enabling parallel updates of multiple tokens rather than the token-by-token process of autoregressive methods, thereby reducing the number of forward passes. Additionally, their bidirectional nature allows them to leverage both past and future context, in contrast to the causal masking that constrains autoregressive models. Early frameworks like D3PM Austin et al. (2021) adapted continuous diffusion to discrete tokens by assigning probability that a token corrupts by either turning into a mask token or random token.
Recent work on score-based discrete diffusion has further advanced the field by providing analytical solutions for the denoising process. Instead of directly modeling transition probabilities, SEDD Lou et al. (2024) uses a score-based approach that learns the gradient of the log-probability and introduces an entropy-based loss function, narrowing the gap to autoregressive models.
@@ 2.3 Optimal Transport Coupling for Flexible Generation
OT Villani et al. (2009) coupling has been well studied in continuous diffusion. Tong et al. (2024) introduce minibatch couplings—either independent (random) or determined by OT—that regularize training and encourage near-straight trajectories. We adopt the term "coupling" from Tong et al. (2024) to denote a matching between an source and target distribution. By incentivizing straighter paths, minibatch OT coupling yields practical benefits such as faster sampling, a pattern also observed in OT-inspired methods like Rectified Flow (Liu et al., 2023) and Stochastic Interpolants (Albergo and Vanden-Eijnden, 2023).
Our OT coupling differs in scope and granularity: we construct a sample-level coupling within each sequence, aligning token positions between an initial state and a target ordering. We instantiate this coupling via OT (and compare to independent coupling in ablations), which preserves relative order and guides tokens along predictable paths while still allowing the inter-set crossings needed for flexible-length infilling. Unlike minibatch couplings aimed primarily at accelerating continuous diffusion models, our coupling enables joint optimization of content and placement of tokens for discrete text diffusion, retaining parallel and bidirectional conditioning and permitting dynamic adjustment of span lengths and token spacing.
@@ 3 Background: Masked Text Diffusion
Discrete diffusion adapts the continuous nature of diffusion processes to text. The simplest and most performant version of discrete diffusion is masked diffusion Lou et al. (2024); Shi et al. (2024); Sahoo et al. (2024). Rather than adding Gaussian noise to continuous values such as pixels, masked text diffusion assigns a probability of masking tokens throughout the forward diffusion process. For the purposes of this paper, masked diffusion can be seen as a masked language model (like BERT Devlin et al. (2019)) that operates at progressively decreasing masking ratios to generate text. Specifically, our text diffusion process follows Score Entropy Discrete Diffusion (SEDD) Lou et al. (2024), which models a score function over a support of $N$ states (token values). The forward diffusion process is described by a continuous-time Markov chain (CTMC):
$\frac{dp_{t}}{dt}=Q_{t}p_{t}\quad p_{0}\approx p_{\mathrm{data}},$ (1)
where $p_{t}\in\mathbb{R}^{N}$ is the column vector of state probabilities at time $t\in[0,T]$, $Q_{t}\in\mathbb{R}^{N\times N}$ is the (time-dependent) transition matrix (its columns sum to 0), and $p_{\mathrm{data}}\in\mathbb{R}^{N}$ is the empirical token distribution at $t=0$. In masked diffusion we include the mask token among the $N$ states and make it absorbing: each non-mask column has 1-1 on its diagonal and +1+1 in the mask row, while the mask column is all zeros, so columns sum to 0 and probability flows from non-mask tokens into the mask token.
SEDD reverses the text diffusion dynamics by learning the score $s_{\theta}(x)_{y}$ of a transition from token $x$ to token $y$ via an entropy-based loss:
$\mathcal{L}_{\text{tok}}=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0})}\sum_{y\neq x_{t}}Q_{t}(x_{t},y)\,\left(s_{\theta}(x_{t},t)_{y}-\right.\\[-2.0pt] \left.\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})}\log s_{\theta}(x_{t},t)_{y}+K\!\left(\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})}\right)\right)\,dt$ (2)
where $x_{0}$ is the start token, $x_{t}$ is the token at time $t$ under the forward CTMC with marginal $p_{t|0}(\cdot|x_{0})$, $Q_{t}(x_{t},y)$ denotes the (row $x_{t}$, column $y$) entry of $Q_{t}$, and $K(a)=a(\log a-1)$ is a scalar regularizer; Finally, to simulate the reverse diffusion process, we either take Euler steps or use an analytical denoiser Lou et al. (2024).
@@ 4 Approach
Prior discrete diffusion models fix token positions during denoising and therefore require oracle positions to infill masked spans. We introduce DDOT, which jointly denoises discrete token values and continuous token positions so infilled spans can move and change length during generation. This enables flexible-length, flexible-position infilling from arbitrary prompts while preserving the parallel, bidirectional nature of discrete diffusion. Figure 2 summarizes DDOT.
4.1 Token Value Diffusion
We partition token values into three disjoint subsets $x=(x^{p},x^{r},x^{\mathrm{pad}})$, which are described as follows: (i) Prompt ($x^{p}$): these slots are clamped to their ground-truth tokens for all $t$; we never mask or resample them, and they serve purely as conditioning during denoising. (ii) Response ($x^{r}$): these are the real tokens to be infilled. They start at ground truth and follow the masked-diffusion forward process from section 3, progressively transitioning toward [mask] as $t$ increases; the reverse chain denoises them back to tokens. (iii) Pad ($x^{\mathrm{pad}}$): these slots hold the vocabulary token [pad] (distinct from [mask]), are not part of $x^{r}$, and exist only to maintain a fixed length $L$. Under the forward process (section 3), [pad] may corrupt to [mask]; in the reverse process, [mask] on these slots denoises back to [pad].
@@ 4.2 Continuous Position Diffusion
To enable diffusion models to move masked tokens to regions that require infilling, DDOT also diffuses positions. Similar to token values, we partition positions into three disjoint subvectors $z_{t}=(z_{t}^{p},\,z_{t}^{r},\,z_{t}^{\mathrm{pad}})$ for prompt, response, and pad positions. Let $\ell=|x^{p}|+|x^{r}|$ be the number of real tokens, with per-subset counts $\ell^{p}=|x^{p}|$ and $\ell^{r}=|x^{r}|=\ell-\ell^{p}$; there are $L$-$\ell$ pad slots.
The noisy positions are sampled i.i.d. as $z_{T}\sim\mathcal{U}(-1,1)^{L}$ (so $z_{T}^{p},z_{T}^{r},z_{T}^{\mathrm{pad}}$ are slices of a single length-$L$ vector). We define the ground-truth positions for the real (non-pad) tokens as the length-$\ell$ subvector over the union of prompt and response indices, $z_{0}^{p\cup r}\in\mathbb{R}^{\ell}$ (equivalently $z_{0}^{p\cup r}=(z_{0}^{p},z_{0}^{r})$), evenly spaced and scaled to the true length:
$z^{p\cup r}_{0,i}=\frac{\ell}{L}\,\frac{2i-(\ell-1)}{\max\{1,\ell-1\}},\qquad i=0,\dots,\ell-1.$ (3)
so real tokens occupy the sub-interval $\ell/L,\,\ell/L]\subset[-1,1]$.
Linear position paths.
Following Albergo and Vanden-Eijnden (2023) and Liu et al. (2023), positions follow straight, noise-free interpolation:
$z_{t}\;=\;(1-t)\,z_{0}+t\,z_{T}.$ (4)
Pad tokens.
Pad tokens are part of the ground-truth vocabulary (distinct from [mask]). Pad positions remain stationary up to the same global scaling:
$z^{\mathrm{pad}}_{0}\;=\;\frac{\ell}{L}\,z^{\mathrm{pad}}_{T}$ (5)
Together with the $z^{p\cup r}$ construction in Equation 3, this induces the full length-$L$ slot vector $z_{0}=(z_{0}^{p},z_{0}^{r},z_{0}^{\mathrm{pad}})$.
4.3 Positional Optimal Transport Coupling
Our preliminary experiments (Table 4) show that naively diffusing continuous token positions performs poorly. It induces a combinatorial explosion of permutations, weakens supervision from the ground truth, and leads to invalid infilling results. We therefore use sample-level optimal-transport (OT) coupling. In particular, we find an OT coupling independently within each set $z^{p}$ and $z^{r}$. Our sample-level OT coupling has the following benefits:
(a) Reduces the permutation space.
OT yields order-preserving, non-crossing trajectories within each set (prompt or response), drastically shrinking the search over permutations. Concretely, for either set $S\in\{p,r\}$ and indices $1\leq i<j\leq\ell^{S}$ in that set’s ground-truth order, we have $z_{t}^{S,(i)}\leq z_{t}^{S,(j)}$ for all $t\in[0,1]$, so within-set crossings cannot occur (while inter-set prompt–response crossings remain allowed), avoiding the combinatorial explosion from arbitrary permutations.
(b) Preserves supervision signal.
Small word-order changes can flip semantics; The OT coupling between $z_{T}$ and $z_{0}$ preserves relative order throughout diffusion and thus maintains a stable learning signal from the ground-truth sequence rather than forcing the model to infer order from noisy, permuted trajectories.
(c) Prevents invalid permutations.
In infilling tasks, the prompt’s relative order is part of the conditioning; any output that reorders prompt tokens is an invalid infilling result. We feed the prompt tokens at the terminal state $T$ in their given order. With OT coupling, the forward $0\!\to\!T$ paths are order-preserving within the prompt set (non-crossing), so the reverse diffusion returns the same within-set order at $t=0$. Without coupling, trajectories can cross and swap prompt order, yielding invalid infills.
Assignment formulation (per set).
We compute the optimal transport coupling within prompt and non-prompt sets. From the noisy positions $z_{T}\in\mathbb{R}^{L}$, sample $\ell^{p}$ entries to form the prompt slice $z_{T}^{p}$; the remaining $L-\ell^{p}$ entries form the non-prompt slice $z_{T}^{\mathrm{np}}$. We compare ground-truth positions (rescaled to [1,1][-1,1]) directly to these slices.
Prompt OT (balanced). One-to-one match: each of the $\ell^{p}$ prompt positions at $t=0$ is assigned to exactly one sampled prompt position at $t=T$, minimizing total displacement:
$\pi^{p}=\operatorname*{arg\,min}_{\pi\in S_{\ell^{p}}}\sum_{i=1}^{\ell^{p}}\Bigl|\tfrac{L}{\ell}\,z^{p}_{0,i}-z^{p}_{T,\pi(i)}\Bigr|,\quad\tilde{z}_{T}^{p}=z_{T}^{p}[\pi^{p}].$ (6)
Where $S_{\ell^{p}}$ is the set of permutations of $\ell^{p}]$ and, for any vector $v$, $v[\pi]$ denotes reindexing by $\pi$: $(v[\pi])_{i}\coloneqq v_{\pi(i)}$.
Response OT (unbalanced). Injective match: the $\ell^{r}$ response positions at $t=0$ are assigned into the non-prompt pool $z_{T}^{\mathrm{np}}$ of size $L-\ell^{p}$, each slot used at most once; unassigned slots become pads:
$\pi^{r}=\operatorname*{arg\,min}_{\pi\in\Pi_{\ell^{r}}^{\,L-\ell^{p}}}\sum_{i=1}^{\ell^{r}}\Bigl|\tfrac{L}{\ell}\,z^{r}_{0,i}-z^{\mathrm{np}}_{T,\pi(i)}\Bigr|,\quad\tilde{z}_{T}^{r}=z_{T}^{\mathrm{np}}[\pi^{r}].$ (7)
Here $S_{\ell^{p}}$ is the set of all permutations of $\ell^{p}]$, and $\Pi_{\ell^{r}}^{\,L-\ell^{p}}$ is the set of injective maps $\ell^{r}]\hookrightarrow[L-\ell^{p}]$.
Pads (leftover positions). The $L$-$\ell$ non-prompt positions not selected by $\pi^{r}$ constitute $z_{T}^{\mathrm{pad}}$. Pads are excluded from OT and follow the stationary path in Equation 5.
After solving the assignments, we reorder the sampled positions and use the reordered slices for interpolation:
$z_{t}^{S}\;=\;(1-t)\,z_{0}^{S}+t\,\tilde{z}_{T}^{S},\qquad S\in\{p,r\}.$ (8)
This per-set OT coupling induces order-preserving, non-crossing trajectories within prompt and response (by construction of $\tilde{z}_{T}^{S}$ and linear paths), while allowing inter-set crossings between prompt and response. See subsection A.4 for visualizations.
Computation.
All couplings are computed per sequence in one dimension, using absolute distance between rescaled ground-truth positions and sampled positions. This choice lets us avoid forming dense cost matrices that are typically used when calculating OT and enables simple, fast implementations that scale favorably in the sequence length.
Prompt OT computation. For the prompt set we must match $\ell^{p}$ ground-truth positions $z_{0}^{p}$ to $\ell^{p}$ sampled positions $z_{T}^{p}$ in a one-to-one manner. In 1D with convex costs, the optimal assignment is equivalent to sorting both vectors and pairing by index. Therefore, we sort $z_{0}^{p}$ and $z_{T}^{p}$ once and obtain the permutation $\pi^{p}$ that maps each sorted index in $z_{0}^{p}$ to the corresponding rank in $z_{T}^{p}$. This procedure achieves the exact minimum from Equation 6, runs in $O(\ell^{p}\log\ell^{p})$ time, and does not materialize a cost matrix.
Response OT computation. We sort $a=z_{0}^{r}$ and $b=z_{T}^{\mathrm{np}}$ in ascending order, then compute the injective, order-preserving match with a standard global-alignment dynamic programming solution (Needleman and Wunsch, 1970). This exactly minimizes Equation 7 and runs in $O(\ell^{r}(L-\ell^{p}))$ time and $O(L-\ell^{p})$ memory.
Complexity and practical overhead. The total per-sequence cost is $O(\ell^{p}\log\ell^{p}+\ell^{r}(L-\ell^{p}))$ time and $O(L)$ memory. For the context sizes considered ($L\leq 1024$), the coupling time is negligible compared to a single denoiser forward pass. In practice we precompute several OT couplings in advance on the CPU and stream the dataset to the GPU, leading to negligible overhead, as seen in subsection 5.3
4.4 Training Objective
At each update we draw a time $t\sim\mathcal{U}(0,1)$, sample noisy positions $z_{T}\sim\mathcal{U}(-1,1)^{L}$, and form the matched targets $\tilde{z}_{T}=(\tilde{z}_{T}^{p},\tilde{z}_{T}^{r},z_{T}^{\mathrm{pad}})$ by solving Equation 6Equation 7 (pads follow Equation 5). We then construct the per-set paths (Equation 8). Independently, for token values we follow SEDD’s forward process (Sec. 3) to obtain $x_{t}$ from $x_{0}$ at the same $t$.
Position loss.
We supervise the position head (Figure 2) on all $L$ slots with a simple MSE toward the straight-path target: